# Theory, and Design Guidelines

### v1

@@ V-B Quantitative Comparison
$\min_{{\boldsymbol{L}},{\boldsymbol{R}},{\boldsymbol{Y}}}\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{L}},{\boldsymbol{R}},{\boldsymbol{Y}};{\boldsymbol{X}})\triangleq\frac{1}{2}\|{\boldsymbol{L}}{\boldsymbol{R}}^{\top}+{\boldsymbol{\Psi}}{\boldsymbol{Y}}-{\boldsymbol{X}}\|_{F}^{2},$ (23)
$\quad\text{s.t. }\|{\boldsymbol{Y}}\|_{0}\leq\kappa.$
The objective parameters ${\boldsymbol{\theta}}^{\rm o}$ of the optimization objective in (23) include the transformation $\boldsymbol{\Psi}$}, as well as the rank and sparsity limits $r,\kappa$.
@@
where $\lambda>0$ controls the sparsity level. Under this relaxation, the update of $\boldsymbol{Y}$} becomes a proximal step for the $\ell_{1}$ term, yielding the soft-thresholding rule. Thus, at iteration $k$, the sparse component is updated as
@@
${\boldsymbol{L}}^{(k+1)}$ $={\boldsymbol{L}}^{(k)}-\eta_{L}\cdot\nabla_{{\boldsymbol{L}}}\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{L}}^{(k)},{\boldsymbol{R}}^{(k)},{\boldsymbol{Y}}^{(k)};{\boldsymbol{X}})\left(({\boldsymbol{R}}^{(k)})^{\top}{\boldsymbol{R}}^{(k)}\right)^{-1},$ (26)
${\boldsymbol{R}}^{(k+1)}$ $={\boldsymbol{R}}^{(k)}-\eta_{R}\cdot\nabla_{{\boldsymbol{R}}}\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{L}}^{(k)},{\boldsymbol{R}}^{(k)},{\boldsymbol{Y}}^{(k)};{\boldsymbol{X}})\left(({\boldsymbol{L}}^{(k+1)})^{\top}{\boldsymbol{L}}^{(k+1)}\right)^{-1},$ (27)
@@
To enable a systematic comparison among the four deep unfolding methodologies, we construct four unfolded variants of the iterative RPCA optimizer described above, each corresponding to one of the methodologies detailed in Section III. All unfolded architectures are implemented with a fixed depth of $K=10$ iterations, and differ solely in the nature of their trainable parameters and the form of their iterative mappings.
@@
    Learning Hyperparameters. The first variant unfolds the iterative RPCA solver with iteration-specific step sizes and thresholds. This approach follows the unfolded RPCA design proposed in [58], in which the hyperparameters of each iteration $k$, denoted ${\boldsymbol{\theta}}^{\rm h}_{k}=\{\eta_{L}^{(k)},\eta_{R}^{(k)},\zeta^{(k)}\}$, are treated as trainable parameters learned from data.
@@
    Learning Objective Parameters. In the second variant, we extend the learning scope to include the parameters of the optimization objective itself. Specifically, the transform $\boldsymbol{\Psi}$} defining the sparsity domain in (23) is allowed to vary across iterations and to be learned from data. Accordingly, the parameters of the $k$th iteration are ${\boldsymbol{\theta}}_{k}=\{{\boldsymbol{\Psi}}_{k},{\boldsymbol{\theta}}^{\rm h}_{k}\}$.
@@
    DNN Inductive Bias. The final methodology departs most significantly from the original optimizer by replacing its gradient-based updates with learned neural mappings. The resulting architecture is based on the one used for learning objective parameters, while having the transformation $\boldsymbol{\Psi}$} be implemented using a two layer convolutional neural network that takes as input the current estimates ${\boldsymbol{L}}^{(k)},{\boldsymbol{R}}^{(k)},{\boldsymbol{Y}}^{(k)}$ and the observation $\boldsymbol{X}$}.
@@
We use the simulation scenario detailed in [58], in which the data matrix sizes are $n_{1}=n_{2}=1000$, the low-rank component has $r=5$, and $10\%$ of the entries of the sparse component are non-zero. All unfolded architectures are trained using the same dataset of RPCA observations $\{{\boldsymbol{X}}_{i}\}$, and are optimized via a supervised loss (6) based on the Frobenius norm deviation in the low rank component. Training is carried out in two stages, corresponding to the schemes outlined in Section III. First, sequential training (L3) is applied, where each iteration is trained independently to minimize the objective over its corresponding output while the preceding iterations are fixed. Subsequently, the entire unfolded model is fine-tuned through end-to-end training (L1), allowing global adjustment of all iteration-specific parameters based on the final reconstruction loss. For the model-based optimizer, the hyperparameters ${\boldsymbol{\theta}}^{\rm h}$ are tuned manually to optimize the loss achieved with sufficient iterations (which here required over $10^{5}$ iterations to converge).
The resulting loss values versus iteration of the unfolded optimizers compared to the original model-based optimizer are reported in Fig. 5. Specifically, in Fig. 5(a) we see that when provided with an accurate description of the optimization (i.e., the same ${\boldsymbol{\theta}}^{\rm o}$ as the one used for generating the data), all unfolded optimizers achieve roughly the same performance, and their gain over the model-based optimizer is in their substantial improvement in latency. It is noted though that DNN-aided methodologies of learned correction term and DNN inductive bias both involve higher parameterization and more complex training. However, when the algorithms are provided with a surrogate description of the inference task, modeled as being provided with a noisy version of ${\boldsymbol{\theta}}^{\rm o}$, we observe in Fig. 5(b) that different unfolding methods lead to different performance levels. Specifically, Fig. 5(b) that the unfolded architecture that dos not rely on explicit parameterization of the objective, i.e., the unfolded model with learned objective, notably outperforms. It is also noted that the DNN inductive bias model, which only uses ${\boldsymbol{\theta}}^{\rm o}$ in its initialization, struggles in learning to deviate substantially from the mismatched objective, and achieves similar results to the unfolded models that use ${\boldsymbol{\theta}}^{\rm o}$ in their processing. Still, all unfolded models notably outperform the model-based optimizer which struggles in terms of both latency and in dealing with mismatched objective.
@@ VI Future Research Directions
A key aspect of deep unfolding lies in the potential to learn per-iteration surrogate objectives that deviate from the original mathematical formulation. Understanding when and why such modifications improve performance, and how they relate to the true task objectives, remains a rich topic for exploration. Developing theoretical and empirical tools for analyzing learned surrogate functions could lead to new principles for balancing interpretability and data-driven abstraction, guiding the design of optimizers that adapt their objectives intelligently throughout the inference process.
@@
Deep unfolding is particularly promising for complex environments where optimization and learning are intertwined across multiple interacting entities. Extending unfolding frameworks to distributed systems [59], where local agents cooperatively optimize shared objectives under communication constraints, has only been explored in very specific settings to-date, e.g., [16]. Similarly, integrating deep unfolding within closed-loop systems [60], such as adaptive control and feedback-based signal processing, opens new opportunities for data-driven design of optimization-based control policies that remain interpretable and verifiable.
@@
Finally, extending deep unfolding to handle dynamic or non-stationary environments is an emerging and practically vital research direction. Incorporating ideas from continual learning, online adaptation, and robust training could yield unfolded optimizers that adapt seamlessly to changing data distributions and system conditions. Such advances would be especially impactful in resource-constrained and mission-critical applications, where the ability to learn and adjust on-the-fly is essential.

### v2

@@ V-B Quantitative Comparison
$\min_{{\boldsymbol{L}},{\boldsymbol{R}},{\boldsymbol{Y}}}\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{L}},{\boldsymbol{R}},{\boldsymbol{Y}};{\boldsymbol{X}})\triangleq\frac{1}{2}\|{\boldsymbol{L}}{\boldsymbol{R}}^{\top}+{\boldsymbol{\Psi}}{\boldsymbol{Y}}-{\boldsymbol{X}}\|_{F}^{2},\quad\text{s.t. }\|{\boldsymbol{Y}}\|_{0}\leq\kappa.$ (23)
The objective parameters ${\boldsymbol{\theta}}^{\rm o}$ of (23) include the transformation $\boldsymbol{\Psi}$}, as well as the rank and sparsity limits $r,\kappa$.
@@
where $\lambda>0$ controls the sparsity level. Under this relaxation, the update of $\boldsymbol{Y}$} becomes a proximal step for the $\ell_{1}$ term, yielding the soft-thresholding rule. Thus, under the assumption that $\boldsymbol{\Psi}$} is an orthonormal transform (i.e., ${\boldsymbol{\Psi}}^{\top}{\boldsymbol{\Psi}}=\mathbf{I}$), at iteration $k$, the sparse component is updated as
@@
${\boldsymbol{L}}^{(k+1)}$ $={\boldsymbol{L}}^{(k)}-\eta_{L}\cdot\nabla_{{\boldsymbol{L}}}\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{L}}^{(k)},{\boldsymbol{R}}^{(k)},{\boldsymbol{Y}}^{(k+1)};{\boldsymbol{X}})\left(({\boldsymbol{R}}^{(k)})^{\top}{\boldsymbol{R}}^{(k)}\right)^{-1},$ (26)
${\boldsymbol{R}}^{(k+1)}$ $={\boldsymbol{R}}^{(k)}-\eta_{R}\cdot\nabla_{{\boldsymbol{R}}}\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{L}}^{(k)},{\boldsymbol{R}}^{(k)},{\boldsymbol{Y}}^{(k+1)};{\boldsymbol{X}})\left(({\boldsymbol{L}}^{(k+1)})^{\top}{\boldsymbol{L}}^{(k+1)}\right)^{-1},$ (27)
@@
To enable a systematic comparison among the deep unfolding methodologies, we construct four unfolded variants of the iterative RPCA optimizer described above, each corresponding to one of the methodologies detailed in Section III. All unfolded architectures are implemented with $K=10$ iterations, and differ solely in the nature of their trainable parameters and the form of their iterative mappings.
@@
    Learning Hyperparameters. The first variant unfolds the iterative RPCA solver with iteration-specific step sizes and thresholds. This approach follows the design proposed in [68], in which the hyperparameters of each iteration $k$, denoted ${\boldsymbol{\theta}}^{\rm h}_{k}=\{\eta_{L}^{(k)},\eta_{R}^{(k)},\zeta^{(k)}\}$, are treated as trainable parameters.
@@
    Learning Objective Parameters. In the second variant, the transform $\boldsymbol{\Psi}$} in (23) is allowed to vary across iterations and is learned from data. Accordingly, the parameters of the $k$th iteration are ${\boldsymbol{\theta}}_{k}=\{{\boldsymbol{\Psi}}_{k},{\boldsymbol{\theta}}^{\rm h}_{k}\}$.
@@
    DNN Inductive Bias. The final methodology replaces gradient-based updates with learned neural mappings. The resulting architecture is based on the one used for learning objective parameters, while having the transformation $\boldsymbol{\Psi}$} be implemented using a two layer convolutional neural network that takes as input the current estimates ${\boldsymbol{L}}^{(k)},{\boldsymbol{R}}^{(k)},{\boldsymbol{Y}}^{(k)}$ and the observation $\boldsymbol{X}$}.
These methodologies span a wide range of parameterizations: while learning only hyperparameters results in highly parameter-efficient models with only $30$ trainable parameters, incorporating learned correction terms ($3\cdot 10^{6}$ parameters) or DNN-based inductive bias ($10^{7}$ parameters) leads to significantly larger models, reflecting the increased flexibility and abstraction of these approaches.
@@
We use the simulation scenario detailed in [68], in which the data matrix sizes are $n_{1}=n_{2}=1000$, the low-rank component has $r=5$, and $10\%$ of the entries of the sparse component are non-zero. We consider two numerical studies: comparing different unfolded architectures arising from the same optimizer (under full and mismatched knowledge of $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}$), and comparing different training losses of (17) for a single architecture. In both cases, all architectures are trained using the same dataset of RPCA observations $\{{\boldsymbol{X}}_{i}\}$, and are optimized via a supervised loss (6) based on the Frobenius norm deviation in the low rank component, enabling systematic comparison in controlled settings.
We first compare the unfolded architectures. Training is carried out in two stages, corresponding to the schemes outlined in Section III. First, sequential training (L3) is applied, where each iteration is trained independently to minimize the objective over its corresponding output while the preceding iterations are fixed. Subsequently, the entire unfolded model is fine-tuned through end-to-end training (L1), allowing global adjustment of all iteration-specific parameters based on the final reconstruction loss. For the model-based optimizer, the hyperparameters ${\boldsymbol{\theta}}^{\rm h}$ are tuned manually to optimize the loss achieved with sufficient iterations (which here required over $10^{5}$ iterations to converge).
The resulting loss values, i.e., the average Frobenius reconstruction error of the low-rank component, versus iteration of the unfolded optimizers compared to the original model-based optimizer are reported in Fig. 5, depicting a zoomed-in view of the initial $K=10$ iterations.. Specifically, in Fig. 5(a) we see that when provided with an accurate description of the optimization (i.e., the same ${\boldsymbol{\theta}}^{\rm o}$ as the one used for generating the data), all unfolded optimizers achieve roughly the same performance, and their gain over the model-based optimizer is in their substantial improvement in latency. It is noted though that DNN-aided methodologies of learned correction term and DNN inductive bias both involve higher parameterization and more complex training. However, when the algorithms are provided with a noisy version of ${\boldsymbol{\theta}}^{\rm o}$, we observe in Fig. 5(b) that different unfolding methods lead to different performance levels. Specifically, the architectures that dos not rely on explicit parameterization of the objective, i.e., learned objective, notably outperforms. It is also noted that the DNN inductive bias model, which only uses ${\boldsymbol{\theta}}^{\rm o}$ in its initialization, struggles in learning to deviate substantially from the mismatched objective, and is outperformed by learned objective. Still, all unfolded models notably outperform the model-based optimizer which struggles in terms of both latency and in dealing with mismatched objective.
@@
Figure 6: Training losses comparison for learned hyperparameters unfolded optimizers, RPCA setup
Next, we compare the different individual training strategies L1-L3, as well the combined (L3 followed by L1 as used in the previous study) for tuning an unfolded optimizer with learned hyperparameters with fully known $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}$. The results, shown in Fig. 6, highlight the individual contribution of each training component. Specifically, end-to-end training (L1) by itself struggles to achieve high performance due to the difficulty of jointly optimizing the parameters. This limitation is partially alleviated by the multi-iteration loss (L2), which provides more direct supervision to the intermediate layers and thereby improves the learning dynamics. Sequential training (L3) enables a more gradual and balanced learning process, as each iteration is trained in a structured manner before moving to later stages. Still, treating each iteration separately may limit the final achievable performance. Finally, combining sequential training with end-to-end fine-tuning yields the best overall results, since it benefits both from the stability of progressive iterative learning and from the global refinement enabled by joint optimization.
@@ VI Future Research Directions
A key aspect of deep unfolding lies in the potential to learn per-iteration surrogate objectives that deviate from the original mathematical formulation. Understanding when and why such modifications improve performance, and how they relate to the true task objectives, remains a rich topic for exploration. Developing theoretical and empirical tools for analyzing learned surrogate functions could lead to new principles for balancing interpretability and abstraction, guiding the design of optimizers that adapt their objectives intelligently throughout the inference process.
@@
Deep unfolding is particularly promising for complex environments where optimization and learning are intertwined across multiple interacting entities. Extending unfolding frameworks to distributed systems [69], where local agents cooperatively optimize shared objectives under communication constraints, has only been explored in very specific settings to date, e.g., [16]. Similarly, integrating deep unfolding within closed-loop systems [70], such as adaptive control and feedback-based signal processing, opens new opportunities for data-driven design of optimization-based control policies that remain interpretable and verifiable.
@@
Finally, extending deep unfolding to handle dynamic or non-stationary environments is a vital research direction. Incorporating ideas from continual learning, online adaptation, and robust training could yield unfolded optimizers that adapt seamlessly to changing data distributions and system conditions. Such advances would be impactful in resource-constrained and mission-critical applications, where the ability to learn and adjust on-the-fly is essential.